OpenStudy (anonymous):

4/30=e^(-0.1x)...solve for x

OpenStudy (anonymous):

\[0.133=e ^{-0.1x}\] \[\ln (0.133)=-0.1x\] is this correct? then just divide both sides by -0.1 ??

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

well that's what I was about to do too... It looks good to me. just fraction \(\frac{4}{30}=\frac{2}{15}\) would be easy if you'd use fractions...but for decimals is not needed :)

OpenStudy (anonymous):

Thanks guys!

OpenStudy (mertsj):

yw